# Can Agents Fix Agent Issues?

### v2

@@ Abstract
LLM-based agent systems are emerging as a new software paradigm and have been widely adopted across diverse domains such as medicine, robotics, and programming. However, maintaining these systems requires substantial effort, as they are inevitably prone to bugs and continually evolve to meet changing external requirements. Therefore, automatically resolving agent issues (i.e., bug reports or feature requests) is a crucial and challenging task. While recent software engineering (SE) agents (e.g., SWE-agent) have shown promise in addressing issues in traditional software systems, it remains unclear how effectively they can resolve real-world issues in agent systems, which differ significantly from traditional software. To fill this gap, we first manually analyze 201 real-world agent issues and identify common categories of agent issues. We then spend 500 person-hours constructing AgentIssue-Bench, a reproducible benchmark comprising 50 agent issue resolution tasks (each with an executable environment and failure-triggering tests). We further evaluate state-of-the-art SE agents on AgentIssue-Bench and reveal their limited effectiveness (i.e., with only 3.33% - 12.67% resolution rates). These results underscore the unique challenges of maintaining agent systems compared to traditional software, highlighting the need for further research to develop advanced SE agents for resolving agent issues. Data and code are available at https://alfin06.github.io/AgentIssue-Bench-Leaderboard/#/.
@@ 1 Introduction
LLM-based agent systems have seen widespread adoption across diverse domains, such as medicine DBLP:journals/corr/abs-2502-18864 , programming devin ; DBLP:journals/corr/abs-2405-15793 , robotics DBLP:conf/iros/KannanVM24 ; DBLP:journals/corr/abs-2311-07226 , psychology qiu2024interactive , and general-purpose personal assistants crewai ; agixt . Driven by rapid advancements, agent systems are emerging as a new software paradigm, playing an increasingly pervasive role in shaping and supporting the full spectrum of human activities.
As products of human intellectual labor, similar as traditional software systems, agent systems are also inevitably prone to quality issues. Recent studies DBLP:journals/corr/abs-2503-13657 have shown that multi-agent systems exhibit diverse failure modes during operation. Moreover, agent systems are continuously evolving to meet changing external requirements, making their maintenance both crucial and labor-intensive. For instance, by May 2025, the agent system MetaGPT metagpt had accumulated over 800 GitHub issues (an issue is typically a bug report or a feature request), highlighting the substantial maintenance workload associated with agent systems.
@@
To understand issues in agent systems, we first perform an empirical study to analyze and catalog real-world agent issues. In particular, we collect 201 real-world GitHub issues along with developer-committed patches from 16 widely-used agent systems. We further build a taxonomy of agent issues with human annotators via grounded theory, resulting in 6 categories and 20 sub-categories of common agent issues. Our taxonomy reveals that real-world agent systems exhibit a diverse range of issues, many of which possess unique characteristics not typically found in traditional software systems. The findings highlight the large engineering effort for maintaining agent systems, confirming that automated issue resolution for agent systems is a challenging and critical problem.
@@
We further evaluate multiple state-of-the-art SE agents (i.e., Agentless DBLP:journals/corr/abs-2407-01489 , AutoCodeRover DBLP:conf/issta/0002RFR24 , and SWE-agent DBLP:journals/corr/abs-2405-15793 ) with both GPT-4o GPT4o and Claude-3.5-Sonnet claude on AgentIssue-Bench. We find that all of the existing SE agents exhibit limited capabilities in resolving agent issues. For instance, only 3.33% to 12.67% of agent issues are correctly resolved, which is significantly lower than the resolution rates achieved when these SE agents are applied to traditional software (e.g., 23.20% - 50.80% resolution rate sweboard ). We further conduct a qualitative analysis to break down the resolution capabilities of SE agents across different categories. Notably, the majority of resolved issues pertain to utility or dependency issues, while the most of LLM-related issues (e.g., compatibility with LLM providers or LLM operation issues) remain unsolved. Overall, our analysis reveals the limitations of current SE agents in resolving agent issues, underscoring the need for building advanced SE agents tailored to the maintenance of agent systems.
@@ 2.1 LLM-based Agent Systems
LLM-based agent systems are emerging as a new software paradigm, which have been widely applied across various fields (e.g., medicine DBLP:journals/corr/abs-2502-18864 , programming devin ; DBLP:journals/corr/abs-2405-15793 , robotics DBLP:conf/iros/KannanVM24 ; DBLP:journals/corr/abs-2311-07226 , psychology qiu2024interactive , and general-purpose personal assistants crewai ; agixt ) with remarkable abilities. An LLM-based agent system DBLP:journals/chinaf/XiCGHDHZWJZZFWXZWJZLYDW25 ; DBLP:journals/fcsc/WangMFZYZCTCLZWW24 typically consists of: (i) an LLM-controlled brain that decomposes and schedules tasks (i.e., planning) and records the historical behaviors (i.e., memory); (ii) a perception component that receives information from the environment; and (iii) an action component that interacts with the environment by invoking external tools. In addition, single-agent systems can collaborate to form multi-agent systems, which can tackle more complex tasks with better flexibility and effectiveness.
Quality problems in LLM-integrated systems. Given the widespread adoption of LLMs, recent work has been looking into quality problems (e.g., bugs or runtime failures) in LLM-integrated systems. For example, Shao et al. shao2025llms catalog the integration bugs in LLM-integrated systems. Different from their work, our work focuses on a specific category of LLM-integrated systems, i.e., LLM-based agent systems. Along this direction, Cemri et al. DBLP:journals/corr/abs-2503-13657 build a taxonomy of failure modes in multi-agent systems. While their work focuses on runtime failure symptoms by analyzing failure trajectories, our taxonomy centers on agent issue resolution by analyzing both real-world user-reported issues and developer-committed patches. Therefore, our work complements existing efforts by providing a perspective on maintaining agent systems, encompassing a broader scope that includes not only bug fixes but also feature requests. Moreover, our work is further different from existing work by introducing the first reproducible benchmark for agent issue resolution and empirically evaluating state-of-the-art SE agents on their ability to resolve agent issues.
@@ 2.2 Software Engineering Agents
Benchmarking issue resolution capabilities of SE agents. With the rise of SE agents, an increasing number of benchmarks have been developed to evaluate their capabilities in addressing real-world issue resolution tasks. For instance, Jimenez et al. swebench build SWE-bench from GitHub issues of 12 Python libraries. Based on SWE-bench, researchers further propose a series of benchmarks, e.g., SWE-bench Lite swebench , SWE-bench verified sweverified , and SWE-bench Lite-S DBLP:journals/corr/abs-2407-01489 , which are refined versions of SWE-bench with additional quality checking. While the SWE-bench series only includes issues of Python software, Zan et al. DBLP:journals/corr/abs-2408-14354 ; DBLP:journals/corr/abs-2502-12115 further propose SWE-bench Java, an issue resolution benchmark for Java software, and Yang et al. DBLP:conf/iclr/YangJZLYWPMSNY025 build SWE-bench Multimodal, comprising frontend issue resolutions tasks from open-source JavaScript libraries. More recently, OpenAI releases SWE-Lancer Diamond swelancer , an issue resolution benchmark with end-to-end tests for both open-source and commercial Expensify expensify software. While existing benchmarks focus exclusively on issue resolution in traditional software systems, our work introduces the first reproducible benchmark targeting issues in agent systems, an emerging software paradigm with features distinct from traditional software. Using this benchmark, we find that current SE agents are still unable to resolve the majority of issue resolution tasks in agent systems.
@@ 3.1.1 Data Collection
Agent system collection. To select diverse and representative agent systems, we first use the GitHub search API to obtain 50 repositories with keywords “AI agents” by Feb 2025. We then manually go through each repository to keep the ones that are LLM-based agent systems (filter out the unrelated ones like paper lists or tutorials); to focus on agent systems with active maintenance, we only keep the ones with more than 1k stars and 30 issues. In this way, we collect 16 agent systems, such as MetaGPT metagpt , AutoGen autogen , GPT-engineer gpt-engineer , and CrewAI crewai . The full list of our analyzed agent systems is in Appendix A.
@@ 3.1.2 Manual Labeling
Taxonomy construction. We manually catalog the 171 agent issues with grounded theory glaser2017discovery . In particular, three human annotators with extensive software development and machine learning experience apply open coding seaman1999qualitative to annotate each issue based on the issue description and the developer-committed patch. They break down each issue into segments and label them with descriptive codes. Then they organize the open codes into structured categories by merging and linking relevant ones. All the annotators further discuss and review the taxonomy until reaching a consensus.
Taxonomy evaluation. We further evaluate our taxonomy on the remaining 30 agent issues. Two annotators independently label each issue. Their annotation reaches a high agreement ratio (Cohen’s Kappa = 0.849); meanwhile there emerge no new categories in addition to our taxonomy during their annotation. The results suggest the generalizability and reliability of our taxonomy.
@@
Category Sub-category Description Incompati- bility with LLM providers (7.46%) Incompatible dependencies (1.99%) Miss or misuse the libraries of LLM providers. Unsupported models (2.99%) Lack the support of recent LLMs. Incompatible parameters to LLM providers (2.49%) Use undefined parameters or miss parameters of LLM query interfaces. Tool-related issues (18.41%) Tool dependency issues (3.48%) Miss or misuse libraries for running agent-invoked tools. Tool configuration issues (2.49%) Misconfigure the settings of agent-invoked tools. Tool implementation errors (8.46%) Incorrect implementation of self-developed agent-invoked tools. Misuse tool interfaces (3.98%) Incorrect tool invocation due to missing/wrong parameters. Memory-related issues (14.43%) Memory initialization issues (2.49%) Incomplete or inconsistent memory states due to database initialization or workspace resetting issues. Memory content errors (10.95%) Incorrect message attributes, misleading content, or content loss caused by faulty storage logic. Memory dependency issues (1.00%) Incorrect internal module dependencies or external libraries required by memory operations. LLM operation issues (31.84%) Model access misconfiguration (7.46%) Model access errors caused by misconfiguration like incorrect model binding or authentication credentials (e.g., API keys). Token usage misconfiguration (3.48%) LLM token management issues such as incorrect limits or pricing. Incorrect model output handlers (8.46%) Incorrect parsing logic for model output or miss handlers for unexpected model behaviors like empty or exceptional responses. Model dependency issues (2.99%) Missing/incompatible libraries related to model operation such as tokenization or transformer dependency conflicts. Context length issues (4.98%) Truncated outputs caused by exceeding context limits or miscalculating context length. Prompt-related issues (4.48%) Suboptimal prompt content or prompt management issues (e.g., fail to set/update prompts). Workflow issues (6.97%) Abnormal agent workflows like hanging or repeated loops. Utility issues (20.90%) Utility implementation issues (9.45%) Implementation errors in LLM-unrelated components (e.g., UI/Docker/logging). Utility dependency issues (4.48%) Miss/incompatible libraries or circular internal dependencies required by general utilities (e.g., testing or file operations). Utility configuration issues (6.97%) External component misconfiguration (e.g., I/O paths, network settings).
@@ 3.2 Taxonomy
LLM operation issues. A large portion (31.84%) of agent-related issues occur during LLM operation. For example, proper configuration of model access and token usage is critical, and misconfiguration in these areas can disrupt agent functionality. Additionally, many issues stem from incorrect handling of model outputs, including: (i) flawed parsing implementations, or (ii) missing handlers for unexpected model responses. Beyond the suboptimal prompt content (e.g., unclear model instructions), prompt management can also introduce risks: as agent systems often maintain a large and evolving pool of prompts, failures in prompt updates or configuration can result in models being queried with incorrect or outdated instructions.
@@ 4.2 Benchmark Details
Evaluation metrics. To evaluate how a technique tackles the agent issue resolution task, we adopt the following metrics to evaluate the patches output by the technique (i.e., SE agents in our experiments). (i) Localization accuracy: if the generated patch modifies the same location as the developer-committed patch, we consider it to have accurately localized the issue. We then compute the percentage of issues for which the generated patches can achieve accurate localization. (ii) Plausible resolution rate: if the generated patch makes the failure-triggering tests pass after being applied, we consider it to plausibly resolve the issue (i.e., denoted as a plausible patch). We then compute the percentage of issues for which the generated patches are plausible patches. (iii) Correct resolution rate: if the generated plausible patch is further semantically-equivalent to the developer-committed patch, we consider it to correctly resolve the issue (i.e., denoted as a correct patch). In particular, given the insufficiency of tests in practice, it is common 10.1145/3663529.3663776 ; 9034821 ; 10.1145/2884781.2884872 that plausible patches are not necessarily correct patches but are just overfitting to the failure-triggering tests. Therefore, only reporting the plausible resolution rate can overestimate the effectiveness of issue resolution techniques. Following the common practice in the program repair area chatrepair ; alpharepair ; xiaase23 ; knod , we further involve human annotators to manually check whether the plausible patches are semantically equivalent to developer-committed patches. We then compute the percentage of issues for which the generated patches are correct patches.

### v3

@@ Abstract
LLM-based agent systems are emerging as a new software paradigm and have been widely adopted across diverse domains such as medicine, robotics, and programming. However, maintaining these systems requires substantial effort, as they are inevitably prone to bugs and continually evolve to meet changing external requirements. Therefore, automatically resolving agent issues (i.e., bug reports or feature requests) is a crucial and challenging task. While recent software engineering (SE) agents (e.g., SWE-agent) have shown promise in addressing issues in traditional software systems, it remains unclear how effectively they can resolve real-world issues in agent systems, which differ significantly from traditional software. To fill this gap, we first manually analyze 201 real-world agent issues and identify common categories of agent issues. We then spend 500 person-hours constructing AgentIssue-Bench, a reproducible benchmark comprising 50 agent issue resolution tasks (each with an executable environment and failure-triggering tests). We further evaluate state-of-the-art SE agents on AgentIssue-Bench and reveal their limited effectiveness (i.e., with only 0.67% - 4.67% resolution rates). These results underscore the unique challenges of maintaining agent systems compared to traditional software, highlighting the need for further research to develop advanced SE agents for resolving agent issues. Data and code are available at https://github.com/alfin06/AgentIssue-Bench.
@@ 1 Introduction
LLM-based agent systems have seen widespread adoption across diverse domains, such as medicine DBLP:journals/corr/abs-2502-18864 ; corrabs250519562 , programming devin ; DBLP:journals/corr/abs-2405-15793 ; chattester ; transagent , robotics DBLP:conf/iros/KannanVM24 ; DBLP:journals/corr/abs-2311-07226 , psychology qiu2024interactive ; YangWCWPGHS024 , and general-purpose personal assistants crewai ; agixt . Driven by rapid advancements, agent systems are emerging as a new software paradigm, playing an increasingly pervasive role in shaping and supporting the full spectrum of human activities.
As products of human intellectual labor, similar as traditional software systems, agent systems are also inevitably prone to quality issues. Recent work DBLP:journals/corr/abs-2503-13657 has shown that multi-agent systems exhibit diverse failure modes during operation. Moreover, agent systems are continuously evolving to meet changing external requirements, making their maintenance both crucial and labor-intensive. For instance, by May 2025, the agent system MetaGPT metagpt had accumulated over 800 GitHub issues (an issue is typically a bug report or a feature request), highlighting the substantial maintenance workload associated with agent systems.
@@
To understand issues in agent systems, we first perform an empirical study to analyze and catalog real-world agent issues. In particular, we collect 201 real-world GitHub issues along with developer-committed patches from 18 widely-used agent systems. We further build a taxonomy of agent issues with human annotators via grounded theory, resulting in 6 categories and 20 sub-categories of common agent issues. Our taxonomy reveals that real-world agent systems exhibit a diverse range of issues, many of which possess unique characteristics not typically found in traditional software systems. The findings highlight the large engineering effort for maintaining agent systems, confirming that automated issue resolution for agent systems is a challenging and critical problem.
@@
We further evaluate multiple state-of-the-art SE agents (i.e., Agentless DBLP:journals/corr/abs-2407-01489 , AutoCodeRover DBLP:conf/issta/0002RFR24 , and SWE-agent DBLP:journals/corr/abs-2405-15793 ) with both GPT-4o GPT4o and Claude-3.5-Sonnet claude on AgentIssue-Bench. We find that all of the existing SE agents exhibit limited capabilities in resolving agent issues. For instance, only 0.67% to 4.67% of agent issues are correctly resolved, which is significantly lower than the resolution rates achieved when these SE agents are applied to traditional software (e.g., 23.20% - 50.80% resolution rate sweboard ). We further conduct a qualitative analysis to break down the resolution capabilities of SE agents across different categories. Notably, the majority of resolved issues pertain to utility or dependency issues, while the most of LLM-related issues (e.g., compatibility with LLM providers or LLM operation issues) remain unsolved. Overall, our analysis reveals the limitations of current SE agents in resolving agent issues, underscoring the need for building advanced SE agents tailored to the maintenance of agent systems.
@@ 2.1 LLM-based Agent Systems
LLM-based agent systems are emerging as a new software paradigm, which have been widely applied across various fields (e.g., medicine DBLP:journals/corr/abs-2502-18864 ; corrabs250519562 , programming devin ; DBLP:journals/corr/abs-2405-15793 , robotics DBLP:conf/iros/KannanVM24 ; DBLP:journals/corr/abs-2311-07226 , psychology qiu2024interactive ; YangWCWPGHS024 , and general-purpose personal assistants crewai ; agixt ) with remarkable abilities. An LLM-based agent system DBLP:journals/chinaf/XiCGHDHZWJZZFWXZWJZLYDW25 ; DBLP:journals/fcsc/WangMFZYZCTCLZWW24 typically consists of: (i) an LLM-controlled brain that decomposes and schedules tasks (i.e., planning) and records the historical behaviors (i.e., memory); (ii) a perception component that receives information from the environment; and (iii) an action component that interacts with the environment by invoking external tools. In addition, single-agent systems can collaborate to form multi-agent systems, which can tackle more complex tasks with better flexibility and effectiveness.
Quality problems in LLM-integrated systems. Given the widespread adoption of LLMs, recent work has been looking into quality problems (e.g., bugs or runtime failures) in LLM-integrated systems. For example, Shao et al. shao2025llms catalog the integration bugs in LLM and RAG systems. Different from their work, our work specifically focuses on LLM agent systems. This scope distinction leads to notable differences in taxonomies, as our taxonomy is framed from an agent architecture perspective (e.g., featuring broader coverage of tool-related issues, finer-grained categorization of memory issues, API and model binding issues). Along this direction, Cemri et al. DBLP:journals/corr/abs-2503-13657 build a taxonomy of failure modes in multi-agent systems. While their work focuses on runtime failure symptoms by analyzing failure trajectories, our taxonomy centers on agent issue resolution by analyzing both real-world user-reported issues and developer-committed patches. Therefore, our work complements existing efforts by providing a perspective on maintaining agent systems, encompassing a broader scope that includes not only bug fixes but also feature requests. Moreover, our work is further different from existing work by introducing the first reproducible benchmark for agent issue resolution and empirically evaluating state-of-the-art SE agents on their ability to resolve agent issues.
@@ 2.2 Software Engineering Agents
Benchmarking issue resolution capabilities of SE agents. With the rise of SE agents, an increasing number of benchmarks have been developed to evaluate their capabilities in addressing real-world issue resolution tasks. For instance, Jimenez et al. swebench build SWE-bench from GitHub issues of 12 Python libraries. Based on SWE-bench, researchers further propose a series of benchmarks, e.g., SWE-bench Lite swebench , SWE-bench verified sweverified , and SWE-bench Lite-S DBLP:journals/corr/abs-2407-01489 , which are refined versions of SWE-bench with additional quality checking. While the SWE-bench series only includes issues of Python software, Zan et al. DBLP:journals/corr/abs-2408-14354 ; DBLP:journals/corr/abs-2502-12115 further propose SWE-bench Java, an issue resolution benchmark for Java software, and Yang et al. DBLP:conf/iclr/YangJZLYWPMSNY025 build SWE-bench Multimodal, comprising frontend issue resolutions tasks from open-source JavaScript libraries. More recently, OpenAI releases SWE-Lancer Diamond swelancer , an issue resolution benchmark with end-to-end tests for Expensify expensify software. While existing benchmarks focus exclusively on issue resolution in traditional software systems, our work introduces the first reproducible benchmark targeting issues in agent systems, an emerging software paradigm with features distinct from traditional software. Using this benchmark, we find that current SE agents are still unable to resolve the majority of issue resolution tasks in agent systems.
@@ 3.1.1 Data Collection
Agent system collection. To select diverse and representative agent systems, we first use the GitHub search API to obtain 50 repositories with keywords “AI agents” by Feb 2025. We then manually go through each repository to keep the ones that are LLM-based agent systems (filter out the unrelated ones like paper lists or tutorials); to focus on agent systems with active maintenance, we only keep the ones with more than 1k stars and 30 issues. In this way, we collect 18 agent systems, such as MetaGPT metagpt , AutoGen autogen , GPT-engineer gpt-engineer , and CrewAI crewai . The full list of our analyzed agent systems is in Appendix A.
@@ 3.1.2 Manual Labeling
Taxonomy construction. We manually catalog the 171 agent issues with grounded theory glaser2017discovery . In particular, three human annotators with extensive software development and machine learning experience apply open coding ChenCLW0L20 ; icseChenYLCLWL21 ; seaman1999qualitative to annotate each issue based on the issue description and the developer-committed patch. They break down each issue into segments and label them with descriptive codes. Then they organize the open codes into structured categories by merging and linking relevant ones. All the annotators further discuss and review the taxonomy until reaching a consensus.
Taxonomy evaluation. We further evaluate our taxonomy on the remaining 30 agent issues. Two annotators independently label each issue. Their annotation reaches a high agreement ratio (Cohen’s Kappa = 0.849); meanwhile there emerge no new categories in addition to our taxonomy during their annotation.
@@
Category Sub-category Description Incompati- bility with LLM providers (7.46%) Incompatible dependencies (1.49%) Miss or misuse the libraries of LLM providers. Unsupported models (2.99%) Lack the support of recent LLMs. Incompatible parameters to LLM providers (2.99%) Use undefined parameters or miss parameters of LLM query interfaces. Tool-related issues (19.90%) Tool dependency issues (3.48%) Miss or misuse libraries for running agent-invoked tools. Tool configuration issues (3.47%) Misconfigure the settings of agent-invoked tools. Tool implementation errors (8.46%) Incorrect implementation of self-developed agent-invoked tools. Misuse tool interfaces (4.48%) Incorrect tool invocation due to missing/wrong parameters. Memory-related issues (14.43%) Memory initialization issues (2.49%) Incomplete or inconsistent memory states due to database initialization or workspace resetting issues. Memory content errors (10.95%) Incorrect message attributes, misleading content, or content loss caused by faulty storage logic. Memory dependency issues (1.00%) Incorrect internal module dependencies or external libraries required by memory operations. LLM operation issues (31.34%) Model access misconfiguration (6.97%) Model access errors caused by misconfiguration like incorrect model binding or authentication credentials (e.g., API keys). Token usage misconfiguration (3.48%) LLM token management issues such as incorrect limits or pricing. Incorrect model output handlers (8.46%) Incorrect parsing logic for model output or miss handlers for unexpected model behaviors like empty or exceptional responses. Model dependency issues (2.99%) Missing/incompatible libraries related to model operation such as tokenization or transformer dependency conflicts. Context length issues (4.98%) Truncated outputs caused by exceeding context limits or miscalculating context length. Prompt-related issues (4.48%) Suboptimal prompt content or prompt management issues (e.g., fail to set/update prompts). Workflow issues (6.47%) Abnormal agent workflows like hanging or repeated loops. Utility issues (20.40%) Utility implementation issues (8.96%) Implementation errors in LLM-unrelated components (e.g., UI/Docker/logging). Utility dependency issues (4.48%) Miss/incompatible libraries or circular internal dependencies required by general utilities (e.g., testing or file operations). Utility configuration issues (6.97%) External component misconfiguration (e.g., I/O paths, network settings).
@@ 3.2 Taxonomy
LLM operation issues. A large portion (31.34%) of agent-related issues occur during LLM operation. For example, proper configuration of model access and token usage is critical, and misconfiguration in these areas can disrupt agent functionality. Additionally, many issues stem from incorrect handling of model outputs, including: (i) flawed parsing implementations, or (ii) missing handlers for unexpected model responses. Beyond the suboptimal prompt content (e.g., unclear model instructions), prompt management can also introduce risks: as agent systems often maintain a large and evolving pool of prompts, failures in prompt updates or configuration can result in models being queried with incorrect or outdated instructions.
@@ 4.2 Benchmark Details
Evaluation metrics. To evaluate how a technique tackles the agent issue resolution task, we adopt the following metrics to evaluate the patches output by the technique (i.e., SE agents in our experiments). (i) Localization accuracy: if the generated patch modifies the same location as the developer-committed patch, we consider it to have accurately localized the issue. We then compute the percentage of issues for which the generated patches can achieve accurate localization. (ii) Plausible resolution rate: if the generated patch makes the failure-triggering tests pass after being applied, we consider it to plausibly resolve the issue (i.e., denoted as a plausible patch). We then compute the percentage of issues for which the generated patches are plausible patches. (iii) Correct resolution rate: if the generated plausible patch is further semantically-equivalent to the developer-committed patch, we consider it to correctly resolve the issue (i.e., denoted as a correct patch). In particular, given the insufficiency of tests in practice, it is common 10.1145/3663529.3663776 ; 9034821 ; 10.1145/2884781.2884872 that plausible patches are not necessarily correct patches but are just overfitting to the failure-triggering tests. Therefore, only reporting the plausible resolution rate can overestimate the effectiveness of issue resolution techniques. Following the common practice in the program repair area chatrepair ; alpharepair ; xiaase23 ; knod , we further involve human annotators to manually check whether the plausible patches are semantically equivalent to developer-committed patches. In particular, two participants independently review each plausible patch by comparing it to the golden patch (i.e., developer-committed), focusing on whether the semantics of the patch fully resolve the underlying issue as intended and do not introduce other functional or semantic errors. If both reviewers agree that the patch is semantically equivalent and correctly resolved the issue, it is labeled as correct. If there is a disagreement between the two reviewers, a third participant would be involved as an adjudicator. The final label is determined only after all three reviewers reach a consensus. We then compute the percentage of issues for which the generated patches are correct patches.